A projectile thrown with an intial velocity u and the angle of projection 15° to the horizontal has a range R. If the same projectile is thrown at an angle of 45° to the horizontal with speed 2u, what will be its range

Grade:10

1 Answers

Vikas TU
7 years ago
When thorwn at 15 degree Range is:
R = u^2sin2thetha/g
R = u^2sin30/g => u^2/2g
or
u^2 = 2gR................(1)
When thrown at 2u speed then,
R’ = 4u^2sin90/g
R’ = 4u^2/g
from eqn (1) we get,
R’ = 8R